OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

(x+6)(x-6) = 0 (x-6)(x-2)

hero (hero):

x^2-36 = 0 x^2 = 36 x = -6,6 x^2-8x + 12 = 0 x^2-6x-2x + 12 = 0 x(x-6)-2(x-6) =0 (x-6)(x-2) = 0 x = 2,6

OpenStudy (anonymous):

hw r they alike ? and diff?

hero (hero):

They're alike because both quadratics. They both have 6 as one of their solutions They're different because they can be solved differently

hero (hero):

Plus the first one doesn't have a "c" value, the second one does